I am having trouble with this one: 8x-(4x+4)=16. x=
Somehow, I came up with 5 but it doesn't look right.

Answer by jorel1380(3719)

8x-(4x+4)=16
8x-4x-4=16
8x-4x=20
4x=20
x=5..